OpenStudy (anonymous):

use elimination to find x and y (a) ((x+2)/3) + ((y-2)/3) = -13 (b) x-y = - 8

OpenStudy (anonymous):

those fractions look like a drag multiply by 3 and start with \[x+2+y-2=-39\]

OpenStudy (anonymous):

or just \[x+y=-39\\ x-y=-8\]

OpenStudy (anonymous):

the y's cancel out and you're left with x=-47 Plug that back into one of the equations (I chose x-y= -8) and you should get y=-39) (-47, -39)

OpenStudy (triciaal):

@svan correct process but watch your basics when you add then 2x = -47